OpenStudy (anonymous):

Help with Algebra II please? Simplify and identify any x-values for which the expression is undefined:

OpenStudy (anonymous):

Simplify and identify any x-values for th which the expression is undefined: 1) x^2 + 3x + 2 / x^2 - 3x - 4 2) 4x^6 / 2x^4 3) x^2 - x^3 / 2x^2 - 5x + 3

OpenStudy (anonymous):

Here is what I got, I would appreciate any feedback / revision! 1) x^2 + 3x + 2 / x^2 - 3x - 4 = (x + 2)(x +1) / (x + 1)(x - 4) == (x + 2) / (x - 4) , x ≠ -1, x ≠ 4 2) 4x6 / 2x4 = 2x^2 / 1 == 2x2 ^^ I know that one is probably wrong. 3) x2 - x3 / 2x2 - 5x + 3 = x2(-x) / (x - 3)(x - 2), x ≠ 3, x ≠ 2

OpenStudy (anonymous):

in the third one when you took x^2 common it will be x^2(1-x) in the numerator not x^2(-x) the first 2 are correct

OpenStudy (anonymous):

and also use ( ) so as the reduce the confusion

OpenStudy (anonymous):

@raunaqrox - Thanks man! I knew something was wrong. So it would be like this; x^2 - x^3 / 2x^2 - 5x + 3 = x^2(1 - x) / (x - 3)(x - 2); x ≠ 3, x ≠ 2 ? After thinking about it, I came up with this solution as well; x^2 - x^3 / 2x2 - 5x + 3 = x^2(1 - x) / 2x(x - 1) - 3(x -1) == x(1- x) / (2x - 3)(x -1) ; x ≠ 3, x ≠ 1. Which one would be correct?

OpenStudy (anonymous):

in the first solution for 2x^2-5x+3 you factorized it to (x-3)(x-2) now lets open (x-3)(x-2) = x^2-5x+6 so this is a mistake

OpenStudy (anonymous):

2nd one is correct

OpenStudy (anonymous):

Alright. Thank you very much for your help!